The following information applies to the questions displayed below. On January 1, 2018, Frontier world issues $39.6 million of 7% bonds, due in 10 years, with interest payable semiannually on June 30 and December 31 each year. The proceeds will be used to build a new ride that combines a roller coaster, a water ride, a dark tunnel, and the great smell of outdoor barbeque, all in one ride 3. Required information 10.00 points Required: 1-a If the market rate is 6%, calculate the issue price. F 0 $ P 0 $1 0 $1, and PV answers in dollars not in millions. Round Market interest rate to 1 decimal place.) 0 $1) Use appropriate factor(s) from the tables provided. Do not round interest rate factors. Enter your Bond Characteristics Face amount Interest payment Market interest rate Periods to maturity Issue price $ 39,600,000 $ 1,584.000 3.0% 1-b. The bonds will issue at Face amount A Premium A Discount
Answer #1

Note: In the question, it is given as 7% Bonds but in 1-a, Interest payment is given as 1584000 in table which denotes 8% Bonds.

Interest payment if 8% coupon rate is taken = 39600000*8%*1/2 = 1584000 i.e. matched with interest payment provided in table of 1-a.

So, we are taking 8% bonds.

3. Calculation of Issue price if market rate is 6%:

1-a. Face amount = 39,600,000

Interest Payment =1,584,000

Market Interest Rate = 3.0%

Periods to maturity = 10*2 = 20

Issue Price = 1584000*Present value annuity factor(3%,20) + 39600000*Present value interest factor(3%,20)

= 1584000*14.8775 + 39600000*0.5537 = 23565960 + 21926520 = 45492480

1-b. The bond will issue at

A Premium

4. Calculation of Issue price if market rate is 7%:

2-a. Face amount = 39,600,000

Interest Payment =1,584,000

Market Interest Rate = 7%/2 = 3.5%

Periods to maturity = 10*2 = 20

Issue Price = 1584000*Present value annuity factor(3.5%,20) + 39600000*Present value interest factor(3.5%,20)

= 1584000*14.2124 + 39600000*0.5026 = 22512441.6 + 19902960 = 42415401.6

2-b. The bond will issue at

A premium

5. Calculation of Issue price if market rate is 8%:

3-a. Face amount = 39,600,000

Interest Payment =1,584,000

Market Interest Rate = 4 %

Periods to maturity = 10*2 = 20

Issue Price = 1584000*Present value annuity factor(4%,20) + 39600000*Present value interest factor(4%,20)

= 1584000*13.59 + 39600000*0.4564 = 21526560 + 18073440 = 39600000

3-b. The bond will issue at

Face amount